Shadow:

One cubic foot holds 7.48 gallons of water, and 1 gallon of water weighs 8.33 pounds. How much does 2.9 cubic feet of water weigh in pounds? In tons?

Shadow:

I know how to do this. Just too lazy to pull out folder paper and/or do it in my head. Also, if any of my classmates need help with this...

Shadow:

\[\frac{ 7.48 gallons }{ ft^{3} } \times \frac{ 8.33pounds }{ galloon } \times 2.9 ft^{3} = 180.69436 pounds\]

Shadow:

\[1 ton = 2000 pounds\]

Shadow:

\[180.69436 pounds \times \frac{ 1 ton }{ 2000 pounds} = 0.09034718 tons \]

Shadow:

For the pounds answer, round to the nearest tenth \[180.7 pounds\] For the tons answer, round to the nearest hundredth \[0.09 tons\]

Shadow:

What is your average speed in miles per hour and in feet per second if you travel a mile in 11 minutes?

Shadow:

\[\frac{ mile }{ 11 minutes } \times \frac{ 60 minutes }{ hr } = 5.454545455 mph \]

Shadow:

Round to the nearest tenth \[5.5mph\]

Shadow:

\[\frac{ mile }{ 11 minutes } \times \frac{ minute }{ 60 seconds } \times \frac{ 5280feet }{ mile } = \frac{ 8ft }{ s }\]

Shadow:

In 1900, a certain country's population was 70,881,132 and its area was 2,811,783 square miles. In 2000, the country's population was 287,906,603 and its area was 3,650,550 square miles. Compute the population densities of both years.

Shadow:

\[PopulationDensity_{1900} = \frac{ 70,881,132 }{ 2,811,783 } = \frac{ 25.20860678 people }{ mile^2 }\]

Shadow:

Round to the nearest whole number \[\frac{ 25 people }{ mi^2 }\]

Shadow:

\[\frac{ 287,906,603 people}{ 3,650,550 mi^2 } = \frac{ 78.86663736 people }{ mi^2 }\]

Shadow:

Round to the nearest whole number, \[\frac{ 79 people }{ mi^2 }\]